OpenStudy (anonymous):

The sum of five times a number and four is equal to negative eleven. What is the number? -2 -3 -4 -5

OpenStudy (anonymous):

is it -2 -3 -4 -5

OpenStudy (jidnyesha):

5n+4=-11 5n=-11-4 5n=-15 n=-3

OpenStudy (anonymous):

thx

OpenStudy (jidnyesha):

ur welcome